# Solvability of fractional semilinear heat equations with distributional forcing terms in Lorentz spaces

### v1

@@ Abstract
Of concern is the Cauchy problem for the fractional semilinear heat equation with a forcing term
@@
where $0<\theta\leqslant 2$, $\gamma>1,\>N\in\mathbb{Z}_{\geqslant 1}$ and $\mu$ is a tempered distribution on $\mathbb{R}^{N}$. In terms of Besov-type spaces, we give necessary conditions and sufficient conditions on $\mu$ for the local-in-time solvability of $\rm{(P)}$ in Lorentz spaces.
@@ 1 Introduction and main results
$\begin{cases}\partial_{t}u+(-\Delta)^{\frac{\theta}{2}}u=|u|^{\gamma-1}u+\mu,&\quad\text{in }\mathbb{R}^{N},\>t>0,\\ u(x,0)\equiv 0,&\quad x\in\mathbb{R}^{N},\end{cases}$ (1.1)
where $\partial_{t}:=\partial/\partial t$, $\gamma>1$, $\theta\in\left]0,2\right]$, $N\in\mathbb{Z}_{\geqslant 1}$ and the forcing term $\mu=\mu(x)$ is a tempered distribution on $\mathbb{R}^{N}$. The symbol $(-\Delta)^{\frac{\theta}{2}}$ represents the fractional power of the Laplace operator $\Delta$ in $\mathbb{R}^{N}$, i.e.,
@@
The aim of the present paper is twofold. First, by utilizing the framework of Besov-type spaces, we construct local-in-time solutions to problem (1.1) for a wide class of forcing terms, including distributions such as the derivative of Dirac delta, as well as Radon measures. Second, in terms of Besov-type spaces, we provide a necessary condition on the forcing term $\mu$ for the existence of local-in-time (possibly) sign-changing solutions belonging to an appropriate function space.
Problem (1.1) has been studied by many mathematicians (see e.g., [3, 9, 11, 12, 13, 14, 27, 28, 29] and references therein) from the perspective of solvability. In these works, attention has been paid to various characteristics of the forcing terms, such as their decay rate at spatial infinity and certain integral quantities. In particular, spatial singularities of forcing terms were considered in [9, 11] and some sufficient conditions and necessary conditions on the forcing terms are demonstrated from the viewpoints of local growth rates and function spaces.
We first focus on sufficient conditions for the existence of solutions to (1.1). Hisa, Ishige and Takahashi [9] aimed to identify the spatially strongest singular profile of $\mu$ for the existence of nonnegative solutions and succeeded in providing some solvable criteria for $\mu$. Their results concerning sufficient conditions can be summarized as follows. There exists a constant $A>0$ depending on $N$, $\theta$ and $\gamma$ such that the following claims hold.
  1. (1)
    Let $\gamma<\frac{N}{N-\theta}$. If a nonnegative Radon measure $\mu$ in $\mathbb{R}^{N}$ satisfies
@@
    for some $T>0$, then problem (1.1) admits a solution in $\mathbb{R}^{N}\times\left[0,T\right[$.
  2. (2)
    Let $\gamma\geqslant\frac{N}{N-\theta}$. If a nonnegative measurable function $\mu:\mathbb{R}^{N}\to\mathbb{R}$ satisfies
    $0\leqslant\mu(x)\leqslant\begin{cases}A|x|^{-\frac{\theta\gamma}{\gamma-1}}+C_{0}&\text{if }\gamma>\frac{N}{N-\theta},\\ A|x|^{-N}\left[\log\left(e+\frac{1}{|x|}\right)\right]^{-\frac{N}{\theta}}+C_{0}&\text{if }\gamma=\frac{N}{N-\theta},\end{cases}$
    for some $C_{0}>0$, then problem (1.1) admits a local-in-time solution.
Theorem (1) is obtained by the contraction mapping theorem in uniformly local $L^{\gamma}$ spaces, while (2) is shown by constructing of some supersolutions.
Ishige, Kawakami and Takada [11] found sufficient conditions on $\mu$ for the case of $\gamma\geqslant\frac{N}{N-\theta}$ by combining the real interpolation method with the uniformly local Zygmund type spaces, which was recently introduced in [10] and applied to Fujita critical Cauchy problem ((1.4) below with $\gamma=1+\theta/N$). Their results were proved by fixed point theorem in uniformly local Lorentz spaces/Zygmund type spaces and were consistent with Theorem (2). For the definition of uniformly local Lorentz spaces, see Definition 1.1 (We do not work with Zygmund type spaces, see [10, 11] for detail). Their sufficient conditions are sharp, but unfortunately, their function spaces does not contain distributional forcing terms.
In this paper, we adopt uniformly local Lorentz-based Besov spaces (see Definition 1.3) to give sufficient conditions on the forcing term $\mu(x)$ for the existence of solutions (see Theorem 1.5). This may be regarded as an inhomogeneous problem version of the works by Kozono and Yamazaki [17] and by Zhanpeisov [30] which constructed local-in-time solutions to problem (1.4) for distributional initial data. In the case of $\gamma>\frac{N}{N-\theta}$, our result covers the previous studies [9, 11] and our result gives a slightly looser condition on $\mu$ than theirs (see Remark 1.6 and Remark 1.7). Our existence theorem also works well in the case of $\gamma<\frac{N}{N-\theta}$; the derivative of Dirac delta is dealt under suitable conditions on $\theta$ and $\gamma$, as well as the sufficient condition in [9] is almost covered (see Remark 1.8).
Secondly, we tackle necessary conditions for the existence of solutions. Hisa, Ishige and Takahashi [9] (building on [8, 13, 14]) found a necessary condition on Radon measure $\mu$ for the existence of nonnegative local-in-time solutions to (1.1). If $u$ is a nonnegative solution to problem (1.1) on $\left]0,T\right[$ for some $T>0$, then the inhomogeneous term $\mu$ is a Radon measure and satisfies the following growth conditions: There exist constants $A_{0}$, $A_{1}>0$ such that the estimate
@@
holds. In addition, when $\gamma=\frac{N}{N-\theta}$, it holds that
@@
These conditions are sharp, considering their sufficient conditions (Claims (1) and (2)). We remark that their argument strongly depends on the nonnegative assumption on $u$ and thus we cannot apply their theory to sign-changing solutions.
Recently, in the study of the Cauchy problem of the Navier–Stokes equation in $\mathbb{R}^{N}$, Kozono, Okada and Shimizu [15, 16] proved that if there exists a Serrin class solution, then the initial data necessarily belongs to a specific homogeneous/inhomogeneous Besov spaces $B^{-1+N/p}_{p,q}$ (They also proved the existence theorem of Serrin class solutions for initial data which belong to $B^{-1+N/p}_{p,q}$). They used the characterization of homogeneous/inhomogeneous Besov spaces via heat semigroup (see e.g., [2, 5, 25]). It is noteworthy that positivity of $u$ is not assumed in their results.
Inspired by their works, we obtain necessary conditions on the forcing term $\mu$ for the existence of solutions to (1.1) in Lorentz spaces (see Theorem 1.9). Although our approach requires solutions to belong to some function spaces, which seems to be a strong assumption compared to [9], we do not need to impose the positivity assumption on $u$. This feature serves as a comparative advantage over [9]. Since the manner in which distributions $\mu$ appear in (1.1) is substantially different from that in (1.4) (cf. (1.2) and (1.5)), a new estimate for the Besov-type norm of $\mu$ needs to be established (see Proposition 2.16).
@@ 1.1 Function spaces and definition of solutions
We introduce some function spaces. Henceforth, the symbol $\left\|f\mid X\right\|$ denotes the norm of $f$ in any normed space $X$.
First we define the spaces for our solutions to problem (1.1). For any mesurable function $f$, we define a map $\alpha_{f}:\mathbb{R}_{>0}\to\left[0,\infty\right]$ as
@@
$f^{\ast}(\lambda)$ $:=\sup\left\{\sigma\in\mathbb{R}_{>0}\mathrel{}\middle|\mathrel{}\alpha_{f}(\sigma)>\lambda\right\}$
, where $|E|$ denotes the $N$-dimensional Lebesgue measure for any measurable set $E$. Here we adopt the convention $\sup\varnothing=0$.
@@
(Lorentz spaces)
@@
    Let $p\in\left]0,\infty\right[$ and $q\in\left]0,\infty\right]$. The uniformly local Lorentz space $L^{p,q}_{\text{ul}}(\mathbb{R}^{N})$ is defined as the set of measurable functions $f:\mathbb{R}^{N}\to\mathbb{R}$ such that
    $\left\|f\mathrel{}\middle|\mathrel{}L^{p,q}_{\mathrm{ul}}\right\|:=\sup_{z\in\mathbb{R}^{N}}\left\|f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p,q}\right\|<\infty,$
    where $\chi_{E}$ is the characteristic function of $E(\subset\mathbb{R})$ and $B(x,r)$ is the open ball in $\mathbb{R}^{N}$ with center $x\in\mathbb{R}^{N}$ and radius $r>0$.
@@
    Lorentz spaces $L^{p,q}$ are just quasi-normed spaces, but for $p\in\left]1,\infty\right[$ and $q\in\left[1,\infty\right]$, they can be equipped with an equivalent norm (see Remark 2.7). In this paper, we only use these spaces in the case of $p>\gamma>1$, $q\geqslant 1$ and thus we regard the spaces $\left(L^{p,q},\left\|\cdot\mathrel{}\middle|\mathrel{}L^{p,q}\right\|\right)$ as Banach spaces by abuse of notation.
@@
    Choosing $p$ and $q$ so that the Lorentz space $L^{p,q}$ is a Banach space, we can prove that the uniformly local Lorentz space $L^{p,q}_{\mathrm{ul}}$ is also a Banach space.
We introduce the inhomogeneous Littlewood–Paley dyadic decomposition in order to define the inhomogeneous Besov-type spaces. One can consult [24] for a general reference. Let $\zeta(t)\in C_{c}^{\infty}\left(\left[0,\infty\right[;\mathbb{R}\right)\>$ satisfy  $0\leqslant\zeta(t)\leqslant 1,\>\zeta(t)\equiv 1$ for $t\leqslant 3/2$ and $\text{supp}\>\zeta\subset\left[0,5/3\right[$. We set
@@
Then, we have $\varphi_{(0)}(\xi),\varphi_{j}(\xi)\in C_{c}^{\infty}(\mathbb{R}^{N})$ for all $j\in\mathbb{Z}$ and
@@
$\Delta_{j}f:=\begin{cases}\mathcal{F}^{-1}\left[\varphi_{(0)}(\xi)\mathcal{F}f\right]&\quad\text{for }j=0,\\ \mathcal{F}^{-1}\left[\varphi_{j}(\xi)\mathcal{F}f\right]&\quad\text{for }j\geqslant 1.\end{cases}$
Now we define the spaces of our forcing terms.
@@
(inhomogeneous Besov–Lorentz spaces) Let $1<p<\infty$, $q,r\in\left[1,\infty\right]$ and $s\in\mathbb{R}$. The inhomogeneous Besov–Lorentz space $B^{s}_{(p,q),r}(\mathbb{R}^{N})$ is defined as the set of tempered distributions $f\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ such that
$\Delta_{j}f$ $\in L^{p,q}_{\mathrm{ul}}(\mathbb{R}^{N})\quad\text{for all }j\in\mathbb{N}\quad\text{and }$
$\left\|f\mid B^{s}_{(p,q),r}\right\|$ $\coloneqq\left\|\left\{2^{sj}\left\|\Delta_{j}f\mid L^{p,q}_{\mathrm{ul}}\right\|\right\}_{j\in\mathbb{N}}\mathrel{}\middle|\mathrel{}l^{r}(\mathbb{N})\right\|<\infty.$
This kind of generalization of Besov spaces was introduced by [17] and has been effectively used in the study of semilinear parabolic equations (see e.g., [22, 30]). Their general properties can be found in [19]. The reason we do not employ Morrey spaces but instead employ Lorentz spaces as a base of Besov spaces lies in the Serrin supercritical case (see Lemma 3.3). We mainly use spaces $B^{s}_{(p,q),r}$ for the case of $q$, $r\in\left\{1,\infty\right\}$. Next we put $S(t)f:=\mathcal{F}^{-1}\left[\exp\left(-t|\xi|^{\theta}\right)\mathcal{F}f\right]$ for every $f\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ and $t>0$ to formulate our definition of solutions.
@@
Let $1<a<\infty$, $s\in\mathbb{R}$, $q,r\in\left[1,\infty\right]$ and $T>0$. For $\mu\in B^{s}_{(a,q),r}(\mathbb{R}^{N})$, we say that a measurable function $u:\mathbb{R}^{N}\times\left]0,T\right[\to\mathbb{R}$ is a solution to (1.1) on $\left]0,T\right[$ if $u$ satisfies
$u\in L^{\infty}(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N}))$
for some $1<p<\infty$ and the following integral equation
@@
There is a possibility that the second term of the right hand side of (1.2) is not a function if $\mu$ is a general tempered distribution. In our existence theorem (Theorem 1.5), such cases are eliminated and it certainly exists as an element of $L^{\infty}\left(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N})\right)$. We do not consider about initial condition in Definition 1.4 but we make a comment on it (see Remark 3.5).

### v2

# Solvability of inhomogeneous fractional semilinear heat equations in Lorentz–Morrey spaces

@@ Abstract
Of concern is the Cauchy problem for the fractional semilinear heat equation with inhomogeneous terms
@@
where $\theta>0$, $\gamma>1,\>N\in\mathbb{Z}_{\geqslant 1}$ and $\mu$ is a tempered distribution on $\mathbb{R}^{N}$. By introducing the Lorentz–Morrey spaces, we overcome limitations of real interpolation in the classical local-Morrey spaces and obtain a sharp integral estimate for the nonlinear term. Moreover, in terms of Besov-type spaces, we give necessary conditions and sufficient conditions on $\mu$ for the local-in-time solvability of $\rm{(P)}$ in Lorentz–Morrey spaces.
@@ 1 Introduction and main results
where $\partial_{t}:=\partial/\partial t$, $\gamma>1$, $\theta>0$, $N\in\mathbb{Z}_{\geqslant 1}$ and the inhomogeneous term $\mu=\mu(x)$ is a tempered distribution on $\mathbb{R}^{N}$. The symbol $(-\Delta)^{\frac{\theta}{2}}$ represents the fractional power of the Laplace operator $\Delta$ in $\mathbb{R}^{N}$, i.e.,
@@
The aim of the present paper is twofold. The first aim is to construct local-in-time solutions to problem (P) for a wide class of inhomogeneous terms, including distributions not previously treated, such as derivatives of the Dirac delta and certain measurable functions that are not locally integrable. Inhomogeneous Besov–Morrey spaces are well-suited to this end; however, due to the insufficiency of real interpolation in the classical Morrey spaces, one cannot expect sharp results. We therefore introduce the Lorentz–Morrey spaces and verify their properties. Second, by establishing a new estimate on Besov-type spaces, we provide a necessary condition on the inhomogeneous term $\mu$ for the existence of local-in-time (possibly) sign-changing solutions belonging to Lorentz–Morrey spaces.
From the perspective of solvability, problem (P) has been studied by many mathematicians (see e.g., [3, 12, 14, 15, 16, 17, 30, 31, 32] and references therein). Among others, spatial singularities of inhomogeneous terms were considered in [12, 14], and sufficient conditions and necessary conditions on the inhomogeneous terms for the local-in-time solvability of (P) have been established in terms of local growth rates and function spaces. Hereafter, we write
$\gamma_{\ast}:=\frac{N}{N-\theta}\,\,\text{ if }\,\,\theta\in\left]0,N\right[\quad\text{and}\quad\gamma_{\ast}:=\infty\,\,\text{ if }\,\,\theta\geqslant N.$
We first review previous work on sufficient conditions for the existence of solutions to (P), and then state our objective. Hisa, Ishige and Takahashi [12] demonstrated some solvable criteria for $\mu$ in the case of $\theta\leqslant 2$. Their results on sufficient conditions can be summarized as follows. There exists a constant $A>0$ depending on $N$, $\theta$ and $\gamma$ such that the following statements hold.
  1. (A)
    Let $\gamma<\gamma_{\ast}$. If a nonnegative Radon measure $\mu$ in $\mathbb{R}^{N}$ satisfies
@@
    for some $T>0$, then problem (P) admits a solution in $\mathbb{R}^{N}\times\left[0,T\right[$.
  2. (B)
    Let $\gamma\geqslant\gamma_{\ast}$. If a nonnegative measurable function $\mu:\mathbb{R}^{N}\to\mathbb{R}$ satisfies
    $\mu(x)\leqslant\begin{cases}A|x|^{-\frac{\theta\gamma}{\gamma-1}}+C_{0}&\text{if }\gamma>\gamma_{\ast},\\ A|x|^{-N}\left[\log\left(e+\frac{1}{|x|}\right)\right]^{-\frac{N}{\theta}}+C_{0}&\text{if }\gamma=\gamma_{\ast},\end{cases}$
    for some $C_{0}>0$, then problem (P) admits a local-in-time solution.
Statement (A) was shown by the contraction mapping theorem in uniformly local $L^{\gamma}$ spaces, while (B) was obtained by constructing of supersolutions.
Ishige, Kawakami and Takada [14] found sufficient conditions on $\mu$ for the case of $\gamma\geqslant\gamma_{\ast}$ and $\theta\leqslant 2$ by combining the real interpolation method with the uniformly local Zygmund type spaces, which was recently introduced in [13] and applied to Fujita critical Cauchy problem ((1.3) below with $\gamma=1+\theta/N$). In light of the statement (B), their sufficient conditions are sharp with respect to the integrability index of function spaces. However, their function spaces do not take the effect of sign-changing of functions into account adequately.
In this paper, we adopt Besov–Lorentz–Morrey (BLM) spaces (see Definition 1.3) as the space of inhomogeneous terms. It is well known that Morrey spaces do not fit well with the real interpolation, but we can recover it by introducing the Lorentz spaces as a base of Morrey spaces (see Proposition 2.8). We note that this is established through a discussion of block spaces in [8]. Here, we give a direct proof, which enables us to obtain more general versions of real interpolation for Morrey-type spaces, including a local Lorentz–Morrey version. As a consequence, we can obtain a sharp sufficient condition for a wide range of inhomogeneous terms (see Theorem 1.5). In the case of $\gamma>\gamma_{\ast}$, our result covers previous studies [12, 14] and takes advantage of Morrey spaces, allowing us to handle a rich class of $\mu$ (see Remark 1.6, 1.7 and 1.8). Our existence theorem also works well in the case of $\gamma<\gamma_{\ast}$; derivatives of the Dirac delta are dealt under suitable conditions on $\theta$ and $\gamma$, as well as the sufficient condition in [12] is covered (see Remark 1.9).
Secondly, we tackle necessary conditions for the existence of solutions. Hisa, Ishige and Takahashi [12] (building on [11, 16, 17]) found a necessary condition on Radon measure $\mu$ for the existence of nonnegative local-in-time solutions to (P). If $u$ is a nonnegative solution to problem (P) on $\left]0,T\right[$ for some $T>0$, then the inhomogeneous term $\mu$ is a Radon measure and satisfies the following local growth conditions: There exist constants $A_{0}$, $A_{1}>0$ such that the estimate
@@
holds. In addition, when $\gamma=\gamma_{\ast}$, it holds that
@@
These conditions are sharp, considering their sufficient conditions (Claims (A) and (B)). We remark that their argument strongly depends on the nonnegative assumption on $u$ and thus we cannot apply their theory to sign-changing solutions.
Recently, in the study of the Cauchy problem of the Navier–Stokes equation in $\mathbb{R}^{N}$, Kozono, Okada and Shimizu [18, 19] proved that if there exists a Serrin class solution, then the initial data necessarily belongs to a specific homogeneous/inhomogeneous Besov spaces $B^{-1+N/p,q}_{p}$ (moreover they proved the existence of Serrin class solutions for initial data which belong to $B^{-1+N/p,q}_{p}$). The key tool of their study is the characterization of homogeneous/inhomogeneous Besov spaces via heat semigroup (see e.g., [2, 5, 28]). It is noteworthy that positivity of $u$ is not assumed in their results. See also [7].
Inspired by their works, we obtain necessary conditions on inhomogeneous terms for the existence of local-in-time solutions to (P) in Lorentz–Morrey spaces (see Theorem 1.10). Although our approach requires solutions to belong to specific function spaces, which seems to be a strong assumption compared to [12], the positivity assumption on $u$ is not needed. This feature serves as a comparative advantage over [12]. Since the manner in which distributions $\mu$ appear in (P) is substantially different from that in (1.3) (cf. (1.1) and (1.4)), a new estimate for the Besov-type norm of $\mu$ needs to be established (see Proposition 2.18).
@@ 1.1 Function spaces and definition of solutions
We introduce some function spaces. Henceforth, the symbol $\left\|f\mathrel{}\middle|\mathrel{}X\right\|$ denotes the quasi-norm of $f$ in any quasi-normed space $X$.
First we define the spaces for our solutions to problem (P). For any mesurable function $f$, we define a function $\alpha_{f}:\mathbb{R}_{>0}\to\left[0,\infty\right]$ as
@@
$f^{\ast}(\lambda)$ $:=\sup\left\{\sigma\in\mathbb{R}_{>0}\mathrel{}\middle|\mathrel{}\alpha_{f}(\sigma)>\lambda\right\}.$
Here we adopt the convention $\sup\varnothing=0$, and $|E|$ denotes the $N$-dimensional Lebesgue measure of a measurable set $E$.
@@
(Lorentz spaces and Lorentz–Morrey spaces)
@@
    Let $0<q\leqslant p<\infty$ and $\kappa\in\left]0,\infty\right]$. The (local) Lorentz–Morrey space $M^{p}_{q,\kappa}(\mathbb{R}^{N})$ is defined as the set of measurable functions $f:\mathbb{R}^{N}\to\mathbb{R}$ such that
    $\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{q,\kappa}\right\|$ $:=\sup_{0<R\leqslant 1}\sup_{z\in\mathbb{R}^{N}}R^{\frac{N}{p}-\frac{N}{q}}\left\|f\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{q,\kappa}(\mathbb{R}^{N})\right\|<\infty,$
    where $\chi_{E}$ is the characteristic function of $E(\subset\mathbb{R}^{N})$ and $B(x,r)$ is the open ball in $\mathbb{R}^{N}$ with center $x\in\mathbb{R}^{N}$ and radius $r>0$.
@@
    Lorentz spaces $L^{p,q}$ are just quasi-normed spaces, but for $p\in\left]1,\infty\right[$ and $q\in\left[1,\infty\right]$, they can be equipped with an equivalent norm (see Remark 2.7). Thus, we regard the spaces $\left(L^{p,q},\left\|\cdot\mathrel{}\middle|\mathrel{}L^{p,q}\right\|\right)$ as Banach spaces by abuse of notation, in the case of $p>1$ and $q\geqslant 1$.
@@
    Choosing $q$ and $r$ so that the Lorentz space $L^{q,r}$ is a Banach space, we can prove that the local Lorentz–Morrey space $M^{p}_{q,r}$ is also a Banach space.
  3. 3.
    Fix $0<q\leqslant p<\infty$. Since it follows from the identity $\alpha_{f^{\ast}}=\alpha_{f}$ and the layer-cake representation that $L^{q,q}=L^{q}$, we define the local Morrey spaces by $M^{p}_{q}:=M^{p}_{q,q}$. It is a Banach space if $q\geqslant 1$.
We introduce the inhomogeneous Littlewood–Paley dyadic decomposition in order to define the inhomogeneous Besov-type spaces. Let $\zeta(t)\in C_{c}^{\infty}\left(\left[0,\infty\right[;\mathbb{R}\right)\>$ satisfy  $0\leqslant\zeta(t)\leqslant 1,\>\zeta(t)\equiv 1$ for $t\leqslant 3/2$ and $\text{supp}\>\zeta\subset\left[0,5/3\right[$. We set
@@
Then we have $\varphi_{(0)}(\xi),\varphi_{j}(\xi)\in C_{c}^{\infty}(\mathbb{R}^{N})$ for all $j\in\mathbb{Z}$ and
@@
$\Delta_{j}f:=\begin{cases}\mathcal{F}^{-1}\left[\varphi_{(0)}(\xi)\mathcal{F}f\right]&\text{for }j=0,\\ \mathcal{F}^{-1}\left[\varphi_{j}(\xi)\mathcal{F}f\right]&\text{for }j\geqslant 1.\end{cases}$
Now we define the spaces of inhomogeneous terms.
@@
(inhomogeneous BLM spaces)
Let $1<q\leqslant p<\infty$, $r,\kappa\in\left[1,\infty\right]$ and $s\in\mathbb{R}$. The inhomogeneous Besov–Lorentz–Morrey (BLM) space $B^{s,r}_{p,q,\kappa}(\mathbb{R}^{N})$ is defined as the set of tempered distributions $f\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ such that
$\Delta_{j}f$ $\in M^{p}_{q,\kappa}(\mathbb{R}^{N})\quad\text{for all }j\in\mathbb{N}\quad\text{and }$
$\left\|f\mid B^{s,r}_{p,q,\kappa}\right\|$ $\coloneqq\left\|\left\{2^{sj}\left\|\Delta_{j}f\mid M^{p}_{q,\kappa}\right\|\right\}_{j\in\mathbb{N}}\mathrel{}\middle|\mathrel{}l^{r}(\mathbb{N})\right\|<\infty.$
We also let $B^{s,r}_{p}$ and $B^{s,r}_{p,q}$ denote the usual Besov spaces and Besov–Morrey spaces (defined in the same way as inhomogeneous BLM spaces by replacing $M^{p}_{q,\kappa}$ with $L^{p}$ and $M^{p}_{q}$), respectively. This kind of generalization of Besov spaces was introduced by Kozono and Yamazaki [20] and has been effectively used in the study of semilinear parabolic equations (see e.g., [26, 33]). One can consult [22, 27] for general references. We mainly use spaces $B^{s,r}_{p,q,\kappa}$ for the case of $r\in\left\{1,\infty\right\}$ and $\kappa=\infty$. Next we put $S(t)f:=\mathcal{F}^{-1}\left[\exp\left(-t|\xi|^{\theta}\right)\mathcal{F}f\right]$ for every $f\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ and $t>0$ to formulate our definition of solutions.
@@
Let $1<b\leqslant a<\infty$, $s\in\mathbb{R}$, $q,r,c\in\left[1,\infty\right]$ and $T>0$. For $\mu\in B^{s,r}_{a,b,c}(\mathbb{R}^{N})$, we say that a measurable function $u:\mathbb{R}^{N}\times\left]0,T\right[\to\mathbb{R}$ is a solution to (P) on $\left]0,T\right[$ if $u$ satisfies
$u\in L^{\infty}(0,T;M^{p}_{q,\infty}(\mathbb{R}^{N}))$
for some $1<q\leqslant p<\infty$ and the following integral equation
@@
There is a possibility that the second term of the right hand side of (1.1) is not a function if $\mu$ is a general tempered distribution. In our existence theorem (Theorem 1.5), such cases are eliminated and it certainly exists as an element of $L^{\infty}\left(0,T;M^{p}_{q,\infty}(\mathbb{R}^{N})\right)$. We do not consider about initial condition in Definition 1.4 but we make a comment on it (see Remark 3.5).